OpenStudy (dtan5457):

Identify if these are direct variation, inverse, or neither. xy=1/5 y/x=8 y=2/x 8y=x

OpenStudy (danjs):

Solve the first one for y

OpenStudy (dtan5457):

1/5/y?

OpenStudy (dtan5457):

/x i mean

OpenStudy (danjs):

Direct Variation\[y~~\alpha~~x\] Indirect Variation \[y~~\alpha~~\frac{ 1 }{ x }\]

OpenStudy (danjs):

If Y gets larger as X gets larger, it is a direct variation, If Y decreases as X gets larger, it is indirect variation

OpenStudy (dtan5457):

it just ask for direct or inverse or neither. the first one is direct or inverse? not sure

OpenStudy (danjs):

\[y = \frac{ 1 }{ 5x }\] Y decreases as X becomes larger, Indirect variation

OpenStudy (dtan5457):

is that inverse?

OpenStudy (danjs):

inverse variation, inversely proportional , indirect variaton.. yeah all the same meaning

OpenStudy (dtan5457):

2nd one i believe is direct because y=8x?

OpenStudy (danjs):

yep, which is that then?

OpenStudy (dtan5457):

direct variation?

OpenStudy (danjs):

y becomes larger as x becomes larger, they are directly proportional

OpenStudy (danjs):

y = k*x

OpenStudy (dtan5457):

3rd one is inverse?

OpenStudy (danjs):

yep

OpenStudy (dtan5457):

last one of y=x/8, is that inverse? x is on the top

OpenStudy (dtan5457):

@DanJS

OpenStudy (danjs):

no that is direct \[y = \frac{ 1 }{ 8 }x\] y = k*x k=1/8